# Minimal Subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{S})$

### v2

@@ 3.2. A diversion into CM elliptic curves
First, the argument in Subsection 1.1 applies also to the CM case and shows that if $E/\mathbb{Q}$ is a CM elliptic curve with bad reduction only at $2$, then $\mathrm{im}\,\rho_{E,2}\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a subgroup for which $\det(\mathrm{im}\,\rho_{E,2})=\mathbb{Z}_{2}^{\times}$ and for which every proper closed subgroup comes from the determinant.
However, the argument in the proof of Theorem 1.2 used in a crucial way that the image of the $S$-adic Galois representation had finite level, and this will not be true for CM curves. Is there a finite set of primes $S$ containing at least one odd prime, a number field $K$ with the property that $K\cap\mathbb{Q}(\zeta_{n})=\mathbb{Q}$ for all positive integers $n$ all of whose prime factors are in $S$, and a CM elliptic curve $E/K$ for which every maximal closed subgroup of the $S$-adic image of Galois comes from $\mathbb{Z}_{S}^{\times}$?
@@ 4. Minimal groups are plentiful
The goal of this section is to show that the minimal groups are plentiful inside of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$. To begin, notice that if $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a minimal group, then $G$ must be (topologically) generated by 2 elements of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$. This is because $G$ must have precisely three maximal closed subgroups, each of which is the preimage under $\det:G\to\mathbb{Z}_{2}^{\times}$ of one of the three maximal closed subgroups of $\mathbb{Z}_{2}^{\times}$.
@@
is zero.
@@
For $n=1$, the power series $f(x_{1})$ is a $p$-adic analytic function on $D$, and it is well-known (see for example [14, Section 6.2]) that the zeros of an analytic function are isolated. From this it follows that $f(x_{1})$ has finitely many zeros in $\mathbb{Z}_{2}$ and a finite set has measure zero. [^3] Strassmann’s theorem gives a more explicit upper bound on the number of zeros of a $p$-adic power series.
@@
Let $E$ be the zero set of $f$. By the Fubini theorem, we have
@@ 5.1. The search for genus 0 groups
We now turn our attention to finding all of the genus 0 minimal subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$ up to conjugacy. Our first challenge is to find a finite box that contains all of the minimal groups. According to [2], the largest index of a subgroup of $\operatorname{PSL}_{2}(\mathbb{Z}_{2})$ with genus 00 is $48$. If $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a minimal group with $|\operatorname{GL}_{2}(\mathbb{Z}_{2}):G|=d$, then surjectivity of the determinant gives that $|\operatorname{SL}_{2}(\mathbb{Z}_{2}):G\cap\operatorname{SL}_{2}(\mathbb{Z}_{2})|=d$ and this implies that the image of $G\cap\operatorname{SL}_{2}(\mathbb{Z}_{2})$ in $\operatorname{PSL}_{2}(\mathbb{Z}_{2})$ has index $d$ or $d/2$. It follows that $d\leq 96$.
@@
Searching for subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/128\mathbb{Z})$ yields 28 minimal groups, 8 of which we already listed in Section 1. For the other 20 minimal groups $G$, the modular curve $X_{G}$ is a conic without rational points. None of these $28$ genus 00 minimal groups contain $-I$. In fact, we enumerated all $7652$ minimal groups of level $\leq 128$ and none of these groups contain $-I$. We conjecture that it is impossible for a minimal subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$ to contain $-I$.
@@ 5.2. Models for the corresponding modular curves
Returning to the 8 minimal groups $G$ that we found using pure thought in Section 1, if $E_{1}$ is the universal elliptic curve over $X_{G}$, then 3 other universal elliptic curves can be obtained by taking of twists of $E_{1}$ by $-1$, $2$, and $-2$. The other 4 examples are $2$-isogenous to the curves obtained this way. It turns out that this relationship persists for most of the other genus 00 minimal groups without points defined over $\mathbb{Q}$. That relationship is made explicit below. The vertical maps are twists by 2, the horizontal maps are twists by $-1$, and the maps from front to back are $2$-isogenies.
@@
For each of the families of elliptic curves we provide simplified models for $E_{1}$. The interested user can then take twists and isogenies to recover the remaining models. For each curve we give a model for, we will define the base field as well as provide an $A$ and $B$ from the base field such that a generic elliptic curve with the corresponding image has the form
@@
For three of the four families of curves, the base field they are defined over is $\mathbb{Q}[a,b]/(a^{2}+b^{2}+1)$ while the remaining curve is defined over $\mathbb{Q}(t)$.
@@ 6. Elliptic curves over imaginary quadratic fields with minimal image
Then $E$ has bad reduction only at prime ideals above $2$ in $\mathbb{Q}(a)$. If $n$ is odd and $n\neq 3$, then $\rho_{E,2}(G_{\mathbb{Q}(a)})$ is minimal and has RSZB label 8.24.0.86. If $n=2$, then $\rho_{E,2}(G_{\mathbb{Q}(a)})$ is minimal and has RSZB label 16.384.9.895. If $n=10$, then $\rho_{E,2}(G_{\mathbb{Q}(a)})$ is minimal and has RSZB label 16.384.9.894.
@@
A computation with models of $2$-adic modular curves shows that $y^{2}=x^{3}+2tx^{2}+(t^{2}+1)x$ is the universal elliptic curve with $2$-adic image 4.12.0.12. It follows that for any $n\geq 1$, the $2$-adic image is contained in 4.12.0.12. Note that this level 4 subgroup does not contain $-I$ and hence $-I\not\in\rho_{E,2}(\mathbb{Q}(a))$ for any $n$.
@@
There are four maximal subgroups of 4.12.0.12 with surjective determinant, and they are 4.24.0.9, 4.24.0.10, 8.24.0.85, and 8.24.0.86. Using a model for 8.12.0.40 (generated by 8.24.0.86 and $-I$) we see that the $2$-adic image $y^{2}=x^{3}+2tx^{2}+(t^{2}+1)x$ is contained in 8.24.0.86 if and only if $t^{2}+1=-2u^{2}$. This certainly occurs if $t=a=\sqrt{-(2^{n}+1)}$ with $n$ odd. Moreover, 8.24.0.86 is a minimal group, and so if $\rho_{E,2}(G_{\mathbb{Q}(a)})$ is contained in it and has surjective determinant, it must equal it.
Suppose now that $n=2$. Let $E^{\prime}:y^{2}=x^{3}+2a^{2}x^{2}+a^{2}(a^{2}+1)x$. Then $E^{\prime}:y^{2}=x^{3}-10x^{2}+20x$ is a quadratic twist of $E$ (by an element of $\mathbb{Q}(a)$). We know from the LMFDB that $\rho_{E^{\prime},2}(G_{\mathbb{Q}})$ is 16.96.3.338. This group and all of its index $2$ subgroups contain $-I$. This implies that $\rho_{E^{\prime},2}(G_{\mathbb{Q}(a)})$ contains $-I$, but we know that $\rho_{E,2}(G_{\mathbb{Q}(a)})$ does not. This implies that $\rho_{E,2}(G_{\mathbb{Q}(a)})$ is an index $2$ subgroup of $\rho_{E^{\prime},2}(G_{\mathbb{Q}(a)})$, which in turn implies that $\rho_{E^{\prime},2}(G_{\mathbb{Q}(a)})$ must be a proper subgroup of $\rho_{E^{\prime},2}(G_{\mathbb{Q}})$. There are four index two subgroups with surjective determinant up to conjugacy, and only one of these is contained in 4.6.0.5 (the subgroup generated by 4.12.0.12 and $-I$). This subgroup is 16.192.9.211 and so this must be $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$. The group 16.192.9.211 has several index $2$ subgroup that do not contain $-I$ and only one of these is contained in 4.12.0.12. This subgroup has label 16.384.9.895, and is minimal.
We apply the same process with $n=10$. Let $E^{\prime}:y^{2}=x^{3}+2a^{2}x^{2}+a^{2}(a^{2}+1)x$. Then $E^{\prime}:y^{2}=x^{3}-2050x^{2}+1049600x$ is a quadratic twist of $E$ (by an element of $\mathbb{Q}(a)$). Since $j(E^{\prime})=257^{3}/2^{8}$, one of the exceptional $j$-invariants from the $2$-adic classification, it follows that $\rho_{E^{\prime},2}(G_{\mathbb{Q}})$ is 16.96.3.335. Once again, this subgroup has no index $2$ subgroups that do not contain $-I$, and therefore $\rho_{E^{\prime},2}(G_{\mathbb{Q}(a)})$ must. However, $\rho_{E,2}(G_{\mathbb{Q}(a)})$ does not, and this implies that $\rho_{E,2}(G_{\mathbb{Q}(a)})$ is an index two subgroup of $\rho_{E^{\prime},2}(G_{\mathbb{Q}(a)})$. There are four index two subgroups of 16.96.3.335 with surjective determinant up to conjugacy, and only one of these is contained in 4.6.0.5. This subgroup is 16.192.9.208 and so this must be $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$. The group 16.192.9.208 has two index $2$ subgroups without $-I$ up to conjugacy and the only one which is contained in 4.12.0.12 is the minimal 16.384.9.894, which must equal $\rho_{E,2}(G_{\mathbb{Q}(a)})$. ∎

### v4

@@ 3.2. A diversion into CM elliptic curves
First, at no point in the argument in Subsection 1.1 do we need to exclude CM elliptic curves. Thus the argument applies also to the CM case and shows that if $E/\mathbb{Q}$ is a CM elliptic curve with bad reduction only at $2$, then $\operatorname{Im}\rho_{E,2^{\infty}}\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a subgroup for which $\det(\operatorname{Im}\rho_{E,2^{\infty}})=\mathbb{Z}_{2}^{\times}$ and for which every proper closed subgroup comes from the determinant.
However, the argument in the proof of Theorem 1.3 used in a crucial way that the image of the $S$-adic Galois representation had finite level, and this will not be true for CM curves. This raises the following open question.
Question 3.7.
Is there a finite set of primes $S$ containing at least one odd prime, a number field $K$ with the property that $K\cap\mathbb{Q}(\zeta_{n})=\mathbb{Q}$ for all positive integers $n$ all of whose prime factors are in $S$, and a CM elliptic curve $E/K$ for which every maximal closed subgroup of the $S$-adic image of Galois comes from $\mathbb{Z}_{S}^{\times}$?
@@ 4. Minimal groups are plentiful
The goal of this section is to show that the minimal groups are plentiful inside of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$. We start with a lemma about how minimal groups are generated.
Lemma 4.1.
If $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a minimal group, then $G$ must be (topologically) generated by 2 elements of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$.
Proof.
Suppose $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a minimal group. Then $G$ must have precisely three maximal closed subgroups, each of which is the preimage under $\det:G\to\mathbb{Z}_{2}^{\times}$ of one of the three maximal closed subgroups of $\mathbb{Z}_{2}^{\times}$. ∎
@@
is zero with respect to natural $p$-adic Haar measure.
@@
For $n=1$, the power series $f(x_{1})$ is a $p$-adic analytic function on $D$, and it is well-known (see for example [15, Section 6.2]) that the zeros of an analytic function are isolated. From this it follows that $f(x_{1})$ has finitely many zeros in $\mathbb{Z}_{2}$ and a finite set has measure zero. [^3] Strassmann’s theorem gives a more explicit upper bound on the number of zeros of a $p$-adic power series.
@@
Let $E$ be the zero set of $f$ and let $\chi_{E}$ be the characteristic function of $E$. By the Fubini theorem, we have
@@ 5.1. The search for genus 0 groups
We now turn our attention to finding all of the genus 0 minimal subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$ up to conjugacy. We focus on genus zero curves because they supply infinitely many examples of elliptic curves whose 2-adic representations are minimal. Our first challenge is to find a finite box that contains all of the genus 0 minimal groups. According to [2, 3], the largest index of a subgroup of $\operatorname{PSL}_{2}(\mathbb{Z}_{2})$ with genus 00 is $48$. If $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{2})$ is a minimal group with $|\operatorname{GL}_{2}(\mathbb{Z}_{2}):G|=d$, then surjectivity of the determinant gives that $|\operatorname{SL}_{2}(\mathbb{Z}_{2}):G\cap\operatorname{SL}_{2}(\mathbb{Z}_{2})|=d$ and this implies that the image of $G\cap\operatorname{SL}_{2}(\mathbb{Z}_{2})$ in $\operatorname{PSL}_{2}(\mathbb{Z}_{2})$ has index $d$ or $d/2$. It follows that $d\leq 96$.
@@
Searching for subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/128\mathbb{Z})$ yields 7652 minimal groups, of which 28 have genus 0. Eight of these were already listed in Section 1. For the other 20 minimal groups $G$, the modular curve $X_{G}$ is a conic without rational points. None of these $28$ genus 00 minimal groups contain $-I$. In fact, none of the minimal groups of level $\leq 128$ contain $-I$. We conjecture that it is impossible for a minimal subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$ to contain $-I$.
@@ 5.2. Models for the corresponding modular curves
Returning to the 8 minimal groups $G$ that we found using pure thought in Section 1, if $E_{1}$ is the universal elliptic curve over $X_{G}$, then 3 other universal elliptic curves can be obtained by taking of twists of $E_{1}$ by $-1$, $2$, and $-2$. We remark here that we are guaranteed that if $E$ is an elliptic curve whose 2-adic image is minimal, then the twists of $E$ by $-1$, $2$ and $-2$ will also have minimal image because $\mathbb{Q}(i)$, $\mathbb{Q}(\sqrt{2})$ and $\mathbb{Q}(\sqrt{-2})$ are the quadratic subfields of $\mathbb{Q}(\zeta_{2^{\infty}})\subseteq\mathbb{Q}(E[2^{\infty}]).$ The other 4 examples are $2$-isogenous to the curves obtained this way. It turns out that this relationship persists for most of the other genus 00 minimal groups without points defined over $\mathbb{Q}$. That relationship is made explicit below. The vertical maps are twists by 2, the horizontal maps are twists by $-1$, and the maps from front to back are $2$-isogenies.
@@
For each of the families of elliptic curves we provide simplified models for $E_{1}$. The interested user can then take twists and isogenies to recover the remaining models. See Remark 5.2 for more details. For each curve we give a model for, we will define the base field as well as provide an $A$ and $B$ from the base field such that a generic elliptic curve with the corresponding image has the form
@@
For three of the four families of curves, the base field they are defined over is the field of fractions of $\mathbb{Q}[a,b]/(a^{2}+b^{2}+1)$ while the remaining curve is defined over $\mathbb{Q}(t)$.
Remark 5.1.
We pause of a moment and give a quick summary of what is happening here. What we know is that every elliptic curve (up to isomorphism) over a number field $K$ in which $-1$ is the sum of two squares and image in 16.48.0.25, 39.96.0.2, or 8.24.0.44 can be obtained by choosing $a$ and $b$ in $K$ with $a^{2}+b^{2}=-1$ and plugging those $a$ and $b$ into our formulas. For 32.96.0.1, every curve (up to isomorphism) defined over $\mathbb{Q}$ with this image can be obtained by choosing a $t\in\mathbb{Q}$ and plugging it into our formulas.
@@ B $2^{3}(t^{2}+1)^{8}(t^{8}-8t^{7}-28t^{6}+56t^{5}+70t^{4}-56t^{3}-28t^{2}+8t+1)^{2}$
Remark 5.2.
Each of the elliptic curves above has a unique rational 2-isogeny. The kernel of this isogeny is exactly $\{\mathcal{O},(0,0)\}.$ A classical result tells us that if $E$ is an elliptic curve of the form
$y^{2}=x^{3}+Ax^{2}+Bx,$
and $\varphi\colon E\to E^{\prime}$ is the 2-isogeny with kernel $\{\mathcal{O},(0,0)\}$, then $E^{\prime}$ is given by
$y^{2}=x^{3}-2Ax^{2}+(A^{2}-4B)x.$
Using this and [21, Proposition 5.4], we can easily compute the models for the other elliptic curves in each family.
@@ 6. Elliptic curves over imaginary quadratic fields with minimal image
A computation with models of $2$-adic modular curves in [17] shows that $y^{2}=x^{3}+2tx^{2}+(t^{2}+1)x$ is the universal elliptic curve with $2$-adic image 4.12.0.12. It follows that for any $n\geq 1$, the $2$-adic image is contained in 4.12.0.12. Note that this level 4 subgroup does not contain $-I$ and hence $-I\not\in\rho_{E,2^{\infty}}(\mathbb{Q}(a))$ for any $n$.
@@
There are four maximal subgroups of 4.12.0.12 with surjective determinant, and they are 4.24.0.9, 4.24.0.10, 8.24.0.85, and 8.24.0.86. Using a model for 8.12.0.40 (generated by 8.24.0.86 and $-I$) we see that the $2$-adic image $y^{2}=x^{3}+2tx^{2}+(t^{2}+1)x$ is contained in 8.24.0.86 if and only if $t^{2}+1=-2u^{2}$. This certainly occurs if $t=a=\sqrt{-(2^{n}+1)}$ with $n$ odd. Moreover, 8.24.0.86 is a minimal group, and so if $\rho_{E,2^{\infty}}(G_{\mathbb{Q}(a)})$ is contained in it and has surjective determinant, it must equal it.
Suppose now that $n=2$. Let $E^{\prime}:y^{2}=x^{3}+2a^{2}x^{2}+a^{2}(a^{2}+1)x$. Then $E^{\prime}:y^{2}=x^{3}-10x^{2}+20x$ is a quadratic twist of $E$ (by an element of $\mathbb{Q}(a)$). We know from the LMFDB that $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}})$ is 16.96.3.338. This group and all of its index $2$ subgroups contain $-I$. This implies that $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$ contains $-I$, but we know that $\rho_{E,2^{\infty}}(G_{\mathbb{Q}(a)})$ does not. This implies that $\rho_{E,2^{\infty}}(G_{\mathbb{Q}(a)})$ is an index $2$ subgroup of $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$, which in turn implies that $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$ must be a proper subgroup of $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}})$. There are four index two subgroups with surjective determinant up to conjugacy, and only one of these is contained in 4.6.0.5 (the subgroup generated by 4.12.0.12 and $-I$). This subgroup is 16.192.9.211 and so this must be $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$. The group 16.192.9.211 has several index $2$ subgroups that do not contain $-I$ and only one of these is contained in 4.12.0.12. This subgroup has label 16.384.9.895, and is minimal.
We apply the same process with $n=10$. Let $E^{\prime}:y^{2}=x^{3}+2a^{2}x^{2}+a^{2}(a^{2}+1)x$. Then $E^{\prime}:y^{2}=x^{3}-2050x^{2}+1049600x$ is a quadratic twist of $E$ (by an element of $\mathbb{Q}(a)$). Since $j(E^{\prime})=257^{3}/2^{8}$, one of the exceptional $j$-invariants from the $2$-adic classification, it follows that $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}})$ is 16.96.3.335. Once again, this subgroup has no index $2$ subgroups that do not contain $-I$, and therefore $\rho_{E^{\prime},2}(G_{\mathbb{Q}(a)})$ must. However, $\rho_{E,2^{\infty}}(G_{\mathbb{Q}(a)})$ does not, and this implies that $\rho_{E,2^{\infty}}(G_{\mathbb{Q}(a)})$ is an index two subgroup of $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$. There are four index two subgroups of 16.96.3.335 with surjective determinant up to conjugacy, and only one of these is contained in 4.6.0.5. This subgroup is 16.192.9.208 and so this must be $\rho_{E^{\prime},2^{\infty}}(G_{\mathbb{Q}(a)})$. The group 16.192.9.208 has two index $2$ subgroups without $-I$ up to conjugacy and the only one which is contained in 4.12.0.12 is the minimal 16.384.9.894, which must equal $\rho_{E,2^{\infty}}(G_{\mathbb{Q}(a)})$. ∎